OpenStudy (anonymous):

Given the function f(x)=8Ix−2I, calculate the following values , f(-2)= f(x+1)= f(x^2+2)=

OpenStudy (theviper):

f(-2)=81*(-2)-21 =-162-21 =-183

OpenStudy (theviper):

M i correct???

OpenStudy (anonymous):

no it's not 81 it's an apsolute value of x-2

OpenStudy (theviper):

I m in X. I was trying to help if I can but sorry:(

OpenStudy (theviper):

@apoorvk plz help

OpenStudy (anonymous):

thanks for trying

OpenStudy (theviper):

yw:)))

OpenStudy (theviper):

@dpaInc, @Callisto , @UnkleRhaukus , @Calcmathlete Plz help:)

OpenStudy (anonymous):

i'll do the first one....:)

OpenStudy (anonymous):

oops it's been done...

OpenStudy (anonymous):

wait... the first one i think is incorrect....

OpenStudy (anonymous):

yes yes

OpenStudy (unklerhaukus):

\[f(x)=8|x−2|\] \[f(-2)=8|-2−2|\]\[f(x+1)=8|(x+1)-2|\]\[f(x^2+2)=8|(x^2+2)-2|\]

OpenStudy (unklerhaukus):

\(|z|\) means the absolute value of z ie the magnitude of \(z\) the length without sign

OpenStudy (callisto):

Replace x by the thing(s) in the bracket and simplify....

OpenStudy (anonymous):

\[f(−2)=32\] \[f(x+1)=8*|x-1| \] \[f(x2+2)=8x^2\] See if you get the same answers :)

OpenStudy (anonymous):

thanks yall